### v1

# Imperial College London

Abstract
Analog computing has been recently revived due to its potential for energy-efficient and highly parallel computations. In this paper, we investigate analog computers that linearly process microwave signals, named microwave linear analog computers (MiLACs), and their applications in signal processing for communications. We model a MiLAC as a multiport microwave network with tunable impedance components, which enables the execution of mathematical operations by reconfiguring the microwave network and applying input signals at its ports. We demonstrate that a MiLAC can efficiently compute the linear minimum mean square error (LMMSE) estimator, widely used in multiple-input multiple-output (MIMO) communications beamforming and detection, with remarkably low computational complexity, unachievable through digital computing. Specifically, the LMMSE estimator can be computed with complexity growing with the square of its input size, rather than the cube, with revolutionary applications to gigantic MIMO beamforming and detection.
Index Terms:
Analog computing, linear minimum mean square error (LMMSE) estimator, microwave networks.
I Introduction
Computing devices play a pivotal role in our society, serving as the backbone of the technological advancement. Modern computers involve digital electronic circuits operating on discrete values (binary) to encode and manipulate data. Digital circuits and signals are widespread today because of their numerous benefits over their analog counterparts [1], such as versatility, more precision and noise robustness, flexibility in storing and processing data, and security.
Despite their popularity, digital computers face two key limitations that may hinder their ability to meet the ever-increasing demand for computational power. First, they are characterized by high power consumption, especially due to the presence of power-hungry analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). Second, their speed is limited by the clock of the digital processors and by the rate of ADCs and DACs. To overcome these two limitations, analog computing has been recently revived as it enables energy-efficient and massively parallelized computations [2]. As defined in [2], analog computing refers to manipulating analog signals in real time to perform specific operations. Modern analog computing devices have been proposed in recent years based on optical systems [3, 4], metamaterials and metasurfaces [5, 6], and resistive memory arrays [7, 8].
Analog computers based on optical systems, or optical analog computers, leverage the speed and parallelism of light to process data, represented as optical signals, at an exceptionally high rate [3, 4]. Tasks related to image processing and machine learning inference have been performed through optical analog computers, as reviewed in [9]. Although optical analog computers have demonstrated significant advantages in low-power and ultra-fast processing, they typically rely on large-scale architectures, as their size cannot be reduced below the operational wavelength.
To realize miniaturized analog computers at scales smaller than the operational wavelength, metamaterials have emerged as artificial materials designed to have unusual electromagnetic (EM) properties [5, 6]. Analog computers based on metamaterials have demonstrated the possibility of solving mathematical equations [10], including multiple equations in parallel exploiting input signals at different frequencies [11]. Particular interest has been dedicated to the 2-dimensional counterpart of metamaterials, namely metasurfaces, further reducing the dimensions of the computing devices. Metasurfaces can be engineered to apply specific transfer functions to the spatial and frequency components of the incident wave, including differentiation, integration, and convolution [5, 6]. Metasurfaces are also expected to play a crucial role in future wireless communications due to their ability to perform computations directly in the wave domain [12, 13, 14].
Besides metamaterials and metasurfaces, resistive memory arrays have also emerged as a prominent technology for building modern analog computing devices [7, 8]. Particularly noteworthy are crosspoint memory arrays, in which resistive memory cells are located at the intersections of horizontal and vertical lines in a matrix shape. The values stored in these memory cells can naturally represent the entries of a matrix, enabling the memory array to perform matrix operations, such as matrix-vector multiplications, directly in the analog domain [15, 16]. The parallelism offered by resistive memory arrays in computing matrix-vector multiplications has been used to advance future multiple-input multiple-output (MIMO) communications in [17, 18].
Several communities have developed modern analog computing solutions to accelerate signal processing operations [3]-[11], [15, 16] as well as to foster wireless communications [12, 13, 14, 17, 18]. However, the efforts of these communities are often siloed, with limited inter-disciplinarity, and this fragmentation has led to a lack of a unified model and shared theoretical understanding. In this paper, we aim to fill this gap by focusing on analog computers that linearly process microwave signals. We develop the model of these analog computers by using rigorous microwave theory, characterize their input-output relationship, and explore their application to signal processing for wireless communications.
The contributions of this paper can be summarized as follows. First, we derive a general model for an analog computer that processes microwave signals linearly, i.e., satisfying the superposition principle, which we denote as microwave linear analog computer (MiLAC). A MiLAC can be regarded as a microwave network having multiple ports, where input signals are applied and output signals are read. This is the first effort in providing a theoretical framework unifying previous works [3]-[18]. Second, we characterize the output of a MiLAC as a function of the input signals and the impedance components constituting its microwave network. The obtained expressions involve matrix-matrix products and matrix inversions, which are computationally expensive operations in the digital domain, as their complexity scales cubically with the matrix size. Nevertheless, they can be instantly computed by a MiLAC once the microwave network is properly reconfigured and the input signals are applied. Third, we show that a MiLAC can be used as a computing device to efficiently calculate the linear minimum mean square error (LMMSE) estimator, widely adopted in wireless communications. Compared to a digital computer, a MiLAC can calculate the LMMSE estimator with a reduction in computational complexity by a factor of $2.5\times 10^{4}$.
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\mathbf{a}^{T}$, $\mathbf{a}^{H}$, $[\mathbf{a}]_{i}$, and $\|\mathbf{a}\|_{2}$ refer to the transpose, conjugate transpose, $i$th element, and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, and $[\mathbf{A}]_{i,k}$ refer to the transpose, conjugate transpose, and $(i,k)$th element of a matrix $\mathbf{A}$, respectively. $\mathbb{C}$ denotes the complex number set. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$ denote the identity matrix and the all-zero matrix with dimensions $N\times N$, respectively, and $\mathbf{0}_{M\times N}$ denote the all-zero matrix with dimensions $M\times N$.
II Model of a Microwave Linear Analog Computer
In this section, we introduce and model a class of analog computers, that we denote as MiLAC. We define a microwave analog computer as an analog computer that processes microwave EM signals, i.e., with frequencies from 300 MHz to 300 GHz [19]. In addition, we introduce a linear analog computer as an analog computer that satisfies the superposition principle, i.e., the output caused by multiple inputs is the sum of the outputs caused by each input individually. Thus, a MiLAC is defined as an analog computer that is both microwave and linear.
A MiLAC can be modeled in general as a linear microwave network having multiple terminals, or ports, where input signals are applied and output signals are read. To enable the MiLAC to compute different functions, its microwave network must be reconfigurable, which can be achieved through the use of tunable impedance components, or, equivalently, tunable admittance components. In the following, we model a MiLAC, detailing its four building blocks: microwave network, input signals, output signals, and tunable admittance components.
II-A Microwave Network
The circuit implementing a MiLAC is in general a linear microwave network, following the definition of MiLAC. According to microwave network theory [19, Chapter 4], a linear microwave network having a certain number of ports $P$ can be characterized by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$, which linearly relates the voltages and currents at the $P$ ports through
$\mathbf{i}=\mathbf{Y}\mathbf{v},$ (1)
where $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$ and $\mathbf{i}=[i_{1},\ldots,i_{P}]^{T}\in\mathbb{C}^{P\times 1}$ are the voltage and current vectors at the ports, with $v_{p}\in\mathbb{C}$ and $i_{p}\in\mathbb{C}$ being the voltage and current at the $p$th port, for $p=1,\ldots,P$.
II-B Input Signals
Assume that the MiLAC receives the input on the first $N$ ports of its reconfigurable microwave network, with $N<P$. Note that this comes with no loss of generality, as the ports can always be reordered to ensure the input is applied on the first $N$ ports. In addition, if an input needs to be applied on an internal node of the microwave network, an additional port can be opened, making this model general. The input is applied by $N$ voltage sources with their series impedance $Z_{0}$, e.g., $Z_{0}=50~{}\Omega$, or, equivalently, series admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Denoting as $u_{n}\in\mathbb{C}$ the voltage of the $n$th voltage source, for $n=1,\ldots,N$, the input vector is given by $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$. Following Ohm’s law, the input voltage $u_{n}$ is related to the voltage $v_{n}$ and current $i_{n}$ at the $n$th port by $i_{n}=Y_{0}(u_{n}-v_{n})$, for $n=1,\ldots,N$, giving
$\mathbf{i}_{1}=Y_{0}\left(\mathbf{u}-\mathbf{v}_{1}\right),$ (2)
where we introduced $\mathbf{v}_{1}=[v_{1},\ldots,v_{N}]^{T}\in\mathbb{C}^{N\times 1}$ and $\mathbf{i}_{1}=[i_{1},\ldots,i_{N}]^{T}\in\mathbb{C}^{N\times 1}$ as the vectors of the voltages and currents at the first $N$ ports of the network, respectively.
Figure 1: Representation of a $P$-port MiLAC with input on $N$ ports, where $P=N+M$.
II-C Output Signals
The output of the MiLAC is read on all the $P$ ports of its reconfigurable microwave network, i.e., is given by the vector $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}$. If an output needs to be read on an internal node of the microwave network, an additional port can be added, confirming the generality of our model. Since the number of input ports $N$ is assumed to be lower than the number of total ports $P$, we can write $P=N+M$, where $M$ is the number of ports with no input signal. To read the output voltage on these $M$ ports without causing undesired reflections within the microwave network, these $M$ ports must be perfectly matched to $Z_{0}$, i.e., they are terminated to ground via an impedance $Z_{0}$, or, equivalently, an admittance $Y_{0}=Z_{0}^{-1}$, as shown in Fig. 1. Thus, according to Ohm’s law, the voltage and current at the $(N+m)$th port are related by $i_{N+m}=-Y_{0}v_{N+m}$, for $m=1,\ldots,M$, where the negative sign arises from the direction of the current $i_{N+m}$. This yields
$\mathbf{i}_{2}=-Y_{0}\mathbf{v}_{2},$ (3)
where we introduced $\mathbf{v}_{2}=[v_{N+1},\ldots,v_{N+M}]^{T}\in\mathbb{C}^{M\times 1}$ and $\mathbf{i}_{2}=[i_{N+1},\ldots,i_{N+M}]^{T}\in\mathbb{C}^{M\times 1}$ as the vectors of voltages and currents at the last $M$ ports of the network, respectively.
II-D Tunable Admittance Components
The microwave network of the MiLAC must be implemented with tunable admittance components, such that its admittance matrix $\mathbf{Y}$ is reconfigurable, allowing the computation of different operations in the analog domain. To realize a MiLAC with maximum flexibility, i.e., whose admittance matrix $\mathbf{Y}$ can be reconfigured to any complex matrix with dimension $P\times P$, we can implement its microwave network with $P^{2}$ tunable admittance components connecting each port to ground and to all other ports. Specifically, port $k$ is connected to ground through an admittance $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, and port $k$ is connected to port $i$ through an admittance $Y_{i,k}\in\mathbb{C}$, $\forall i\neq k$. An example of such a MiLAC with maximum flexibility is reported in Fig. 2, where we have $P=4$ ports and $N=2$ input signals.
This topology of tunable admittance components allows us to arbitrarily reconfigure the admittance matrix $\mathbf{Y}$ since its $(i,k)$th element is given by
$\left[\mathbf{Y}\right]_{i,k}=\begin{cases}-Y_{i,k}&i\neq k\\ \sum_{p=1}^{P}Y_{p,k}&i=k\end{cases},$ (4)
for $i,k=1,\ldots,P$, following microwave network theory [19, Chapter 4]. By inverting (4), we can derive the expression of the tunable admittance components as a function of any given admittance matrix $\mathbf{Y}$ as
$Y_{i,k}=\begin{cases}-\left[\mathbf{Y}\right]_{i,k}&i\neq k\\ \sum_{p=1}^{P}\left[\mathbf{Y}\right]_{p,k}&i=k\end{cases},$ (5)
for $i,k=1,\ldots,P$, showing that the $P^{2}$ tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ can be adjusted to allow $\mathbf{Y}$ to assume any arbitrary value.
Figure 2: Representation of a $4$-port MiLAC with input on $2$ ports.
III Analysis of a Microwave Linear Analog Computer
We have introduced and modeled the concept of MiLAC, highlighting that it consists of a reconfigurable microwave network with multiple ports where input signals are applied and output signals are read. In this section, we characterize the output $\mathbf{v}$ as a function of the input $\mathbf{u}$ and the admittance matrix of the microwave network $\mathbf{Y}$.
We begin by noticing that a MiLAC is described by the system of linear equations (1), (2), and (3), given by
$\begin{cases}\mathbf{i}&=\mathbf{Y}\mathbf{v}\\ \mathbf{i}_{1}&=Y_{0}\left(\mathbf{u}-\mathbf{v}_{1}\right)\\ \mathbf{i}_{2}&=-Y_{0}\mathbf{v}_{2}\end{cases},$ (6)
which fully characterizes the MiLAC in Fig. 1. To solve (6), we rewrite it in a more compact form as
$\begin{cases}\mathbf{i}&=\mathbf{Y}\mathbf{v}\\ \mathbf{i}&=Y_{0}\left(\tilde{\mathbf{u}}-\mathbf{v}\right)\end{cases},$ (7)
where we exploited the fact that $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$ and $\mathbf{i}=[\mathbf{i}_{1}^{T},\mathbf{i}_{2}^{T}]^{T}$, and introduced the vector $\tilde{\mathbf{u}}\in\mathbb{C}^{P\times 1}$ as $\tilde{\mathbf{u}}=[\mathbf{u}^{T},\mathbf{0}_{M\times 1}^{T}]^{T}$. By noticing that (7) yields
$\mathbf{Y}\mathbf{v}=Y_{0}\left(\tilde{\mathbf{u}}-\mathbf{v}\right),$ (8)
we can readily express the output $\mathbf{v}$ as
$\mathbf{v}=\mathbf{P}^{-1}\tilde{\mathbf{u}},$ (9)
where the matrix $\mathbf{P}\in\mathbb{C}^{P\times P}$ is defined as
$\mathbf{P}=\frac{\mathbf{Y}}{Y_{0}}+\mathbf{I}_{P},$ (10)
giving the output $\mathbf{v}$ as a function of the admittance matrix $\mathbf{Y}$ and the input $\mathbf{u}$. By denoting the inverse of $\mathbf{P}$ as
$\mathbf{Q}=\mathbf{P}^{-1},$ (11)
partitioned as
$\mathbf{Q}=\begin{bmatrix}\mathbf{Q}_{11}&\mathbf{Q}_{12}\\ \mathbf{Q}_{21}&\mathbf{Q}_{22}\end{bmatrix},$ (12)
with $\mathbf{Q}_{11}\in\mathbb{C}^{N\times N}$, $\mathbf{Q}_{12}\in\mathbb{C}^{N\times M}$, $\mathbf{Q}_{21}\in\mathbb{C}^{M\times N}$, $\mathbf{Q}_{22}\in\mathbb{C}^{M\times M}$, the output vector $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$ given by (9) can be further expressed as
$\begin{bmatrix}\mathbf{v}_{1}\\ \mathbf{v}_{2}\end{bmatrix}=\begin{bmatrix}\mathbf{Q}_{11}\\ \mathbf{Q}_{21}\end{bmatrix}\mathbf{u},$ (13)
solely depending on the blocks $\mathbf{Q}_{11}$ and $\mathbf{Q}_{21}$ and the input $\mathbf{u}$. To explicitly express $\mathbf{Q}_{11}$ and $\mathbf{Q}_{21}$ in (13) as a function of the admittance matrix $\mathbf{Y}$, we exploit the following proposition.
Proposition 1.
Let $\mathbf{P}\in\mathbb{C}^{P\times P}$ be an invertible matrix, and $\mathbf{P}^{-1}\in\mathbb{C}^{P\times P}$ its inverse, partitioned as
$\mathbf{P}=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{bmatrix},\;\mathbf{P}^{-1}=\begin{bmatrix}\mathbf{A}^{\prime}&\mathbf{B}^{\prime}\\ \mathbf{C}^{\prime}&\mathbf{D}^{\prime}\end{bmatrix},$ (14)
where $\mathbf{A},\mathbf{A}^{\prime}\in\mathbb{C}^{N\times N}$, $\mathbf{B},\mathbf{B}^{\prime}\in\mathbb{C}^{N\times M}$, $\mathbf{C},\mathbf{C}^{\prime}\in\mathbb{C}^{M\times N}$, $\mathbf{D},\mathbf{D}^{\prime}\in\mathbb{C}^{M\times M}$, and $P=N+M$. Then, we have the following results:
  1. 1.
    If $\mathbf{A}$ is invertible, it holds
    $\mathbf{A}^{\prime}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\mathbf{B}\left(\mathbf{C}\mathbf{A}^{-1}\mathbf{B}-\mathbf{D}\right)^{-1}\mathbf{C}\mathbf{A}^{-1},$ (15)
    $\mathbf{C}^{\prime}=\left(\mathbf{C}\mathbf{A}^{-1}\mathbf{B}-\mathbf{D}\right)^{-1}\mathbf{C}\mathbf{A}^{-1},$ (16)
  2. 2.
    If $\mathbf{D}$ is invertible, it holds
    $\mathbf{A}^{\prime}=-\left(\mathbf{B}\mathbf{D}^{-1}\mathbf{C}-\mathbf{A}\right)^{-1},$ (17)
    $\mathbf{C}^{\prime}=\mathbf{D}^{-1}\mathbf{C}\left(\mathbf{B}\mathbf{D}^{-1}\mathbf{C}-\mathbf{A}\right)^{-1},$ (18)
  3. 3.
    If $\mathbf{A}$ and $\mathbf{D}$ are both invertible, (15)-(16) and (17)-(18) are equivalent.
Proof.
The proposition follows from [20, Theorem 2.1]. ∎
To rewrite (13) as a function of $\mathbf{Y}$ by using Proposition 1, we partition $\mathbf{P}$ and $\mathbf{Y}$ as
$\mathbf{P}=\begin{bmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\end{bmatrix},\;\mathbf{Y}=\begin{bmatrix}\mathbf{Y}_{11}&\mathbf{Y}_{12}\\ \mathbf{Y}_{21}&\mathbf{Y}_{22}\end{bmatrix},$ (19)
where $\mathbf{P}_{11},\mathbf{Y}_{11}\in\mathbb{C}^{N\times N}$, $\mathbf{P}_{12},\mathbf{Y}_{12}\in\mathbb{C}^{N\times M}$, $\mathbf{P}_{21},\mathbf{Y}_{21}\in\mathbb{C}^{M\times N}$, and $\mathbf{P}_{22},\mathbf{Y}_{22}\in\mathbb{C}^{M\times M}$, giving
$\begin{bmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\end{bmatrix}=\begin{bmatrix}\mathbf{Y}_{11}/Y_{0}+\mathbf{I}_{N}&\mathbf{Y}_{12}/Y_{0}\\ \mathbf{Y}_{21}/Y_{0}&\mathbf{Y}_{22}/Y_{0}+\mathbf{I}_{M}\end{bmatrix},$ (20)
as a consequence of (10). Thus, assuming $\mathbf{P}$ to be invertible, with $\mathbf{P}_{11}$ or $\mathbf{P}_{22}$ invertible, the following results hold as a consequence of (13):
  1. 1.
    If $\mathbf{P}_{11}$ is invertible, we have
    $\mathbf{v}_{1}=\Big{(}\mathbf{P}_{11}^{-1}-\mathbf{P}_{11}^{-1}\mathbf{P}_{12}\Big{.}\\ \left.\times\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u},$ (21)
    $\mathbf{v}_{2}=\left(\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}.$ (22)
  2. 2.
    If $\mathbf{P}_{22}$ is invertible, we have
    $\mathbf{v}_{1}=-\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\mathbf{u},$ (23)
    $\mathbf{v}_{2}=\left(\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\right)\mathbf{u}.$ (24)
  3. 3.
    If $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ are both invertible, (21)-(22) and (23)-(24) are equivalent.
Remarkably, (21)-(22) and (23)-(24) contain the expressions of the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as a function of the input $\mathbf{u}$ and the blocks of $\mathbf{P}$, which are directly related to the blocks of $\mathbf{Y}$ by (20).
Assuming that the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$ are allowed to take any arbitrary complex value, a MiLAC can compute (21)-(22) and (23)-(24) given any blocks $\mathbf{P}_{11}$, $\mathbf{P}_{12}$, $\mathbf{P}_{21}$, and $\mathbf{P}_{22}$. This can be achieved by first computing $\mathbf{Y}$ from any arbitrary $\mathbf{P}$ as
$\mathbf{Y}=Y_{0}\mathbf{P}-Y_{0}\mathbf{I}_{P},$ (25)
which is obtained by inverting (10), and then by setting the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$ as a function of $\mathbf{Y}$ as in (5). By substituting (25) into (5), we obtain that the tunable admittance components can be reconfigured as a function of any arbitrary $\mathbf{P}$ as
$Y_{i,k}=\begin{cases}-Y_{0}\left[\mathbf{P}\right]_{i,k}&i\neq k\\ Y_{0}\sum_{p=1}^{P}\left[\mathbf{P}\right]_{p,k}-Y_{0}&i=k\end{cases},$ (26)
for $i,k=1,\ldots,P$, allowing the MiLAC to compute (21)-(22) and (23)-(24) for any given $\mathbf{P}_{11}$, $\mathbf{P}_{12}$, $\mathbf{P}_{21}$, and $\mathbf{P}_{22}$.
Note that (21)-(22) and (23)-(24) would require computationally expensive matrix-matrix products and matrix inversion operations if computed through a digital computer. However, they can be quickly computed by a MiLAC as the EM signal propagates within the microwave network at light speed. In other words, after properly setting the input $\mathbf{u}$ and the tunable components $\{Y_{i,k}\}_{i,k=1}^{P}$, (21)-(22) and (23)-(24) can be computed with complexity $\mathcal{O}(1)$, independently from the input size $N$ and output size $P=N+M$. Remarkably, the expressions in (22) and (24) are particularly relevant for computing the LMMSE estimator, as it is clarified in Section IV.
TABLE I: Similarity between MiLAC output and LMMSE estimator.
MiLAC output LMMSE estimator
$\mathbf{v}_{2}=\left(\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\left(\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\right)\mathbf{y}$
$\mathbf{v}_{2}=\left(\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\right)\mathbf{u}$ $\hat{\mathbf{x}}_{\textnormal{LMMSE},2}=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}$
IV Analog Computing for the LMMSE Estimator
We have introduced a MiLAC as an analog computer that linearly processes microwave signals and we characterized its output as a function of its input and microwave network admittance matrix. In this section, we show that a MiLAC can be used to compute the LMMSE estimator for linear observation processes with significantly reduced computational complexity.
IV-A LMMSE Estimator for Linear Observation Processes
Consider a linear observation process in which the known random vector $\mathbf{y}\in\mathbb{C}^{Y\times 1}$, also denoted as the observation vector, is given by
$\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n},$ (27)
where $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. According to estimation theory, the LMMSE estimator $\hat{\mathbf{x}}_{\text{LMMSE}}$ of $\mathbf{x}$ is defined as the linear function of $\mathbf{y}$ minimizing the mean square error (MSE) $\text{E}[\|\hat{\mathbf{x}}_{\text{LMMSE}}-\mathbf{x}\|_{2}^{2}]$ [21, Chapter 2]. The expression of the LMMSE estimator for a linear observation process is available in closed form, as given by the following proposition.
Proposition 2.
Considering the observation process in (27), the LMMSE estimator $\hat{\mathbf{x}}_{\text{LMMSE}}$ of $\mathbf{x}$ is equivalently given by
$\hat{\mathbf{x}}_{\textnormal{LMMSE},1}$ $=\left(\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\right)\mathbf{y},$ (28)
$\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$ $=\left(\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\right)\mathbf{y}.$ (29)
Proof.
The proposition follows from [21, Theorem 12.1]. ∎
Interestingly, there is a remarkable similarity between the LMMSE expressions in (28)-(29) and the expressions characterizing the output $\mathbf{v}_{2}$ of a MiLAC in (22) and (24), as highlighted in Tab. I. Thus, we can efficiently compute the LMMSE estimator for linear observation processes with a MiLAC, as specified in the following.
IV-B Computing the LMMSE Estimator with a MiLAC
The expressions of the LMMSE estimator in (28) and (29) can be obtained on the output $\mathbf{v}_{2}$ of a MiLAC, i.e., $\mathbf{v}_{2}=\hat{\mathbf{x}}_{\text{LMMSE},1}$ and $\mathbf{v}_{2}=\hat{\mathbf{x}}_{\text{LMMSE},2}$, by exploiting (22) and (24), respectively. To this end, the input vector of the MiLAC is set as $\mathbf{u}=\mathbf{y}$, and the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ are set according to (26), with $\mathbf{P}=\mathbf{P}_{\text{LMMSE}}$, where
$\mathbf{P}_{\text{LMMSE}}=\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$ (30)
are two values of $\mathbf{P}_{\text{LMMSE}}$ giving the desired output. Note that this MiLAC has $N=Y$ input ports and $M=X$ output ports with no input, giving a total of $P=X+Y$ ports. Interestingly, analog computers have been proposed to compute generalized regression, ridge regression, and linear regression through resistive memory arrays [15, 16, 17], all operations that can be seen as special cases of the LMMSE estimator, confirming the generality of our model.
V Performance Evaluation
A MiLAC can compute the LMMSE estimator with significantly reduced computational complexity compared to a digital computer, as no matrix-matrix product or matrix inversion operation is required. While a MiLAC computes the LMMSE estimator entirely in the analog domain requiring no digital operation, computing the values of its tunable admittance components requires a digital computer. Thus, the computational complexity of computing the LMMSE estimator with a MiLAC is given by the complexity of digitally computing the tunable admittance components in (26). By exploiting the fact that $\mathbf{P}$ in (30) is Hermitian and precomputing offline the computations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$, (26) require approximately $2XY$ and $4XY$ real operations to be digitally computed for the two cases $i\neq k$ and $i=k$, respectively, giving $6XY$ real operations in total. Conversely, the number of real operations required to compute (28) and (29) through digital computing is $38XY^{2}+8X^{2}Y+8X^{3}/$ 3 and $38X^{2}Y+8XY^{2}+8Y^{3}/$ 3, respectively, due to the two matrix-matrix products and the matrix inversion, whose complexity is assessed in the Appendix. As (28) and (29) are equivalent, the complexity of computing the LMMSE estimator digitally is $\mathrm{min}\{8XY^{2}+8X^{2}Y+8X^{3}/3,8X^{2}Y+8XY^{2}+8Y^{3}/3\}$, i.e., $8(XY^{2}+X^{2}Y+\mathrm{min}\{X^{3},Y^{3}\}/3)$. The complexity of computing the LMMSE estimator with a MiLAC and by digital computing are compared in Fig. 3, showing a gain of $2.5\times 10^{4}$ times in the case $X=Y=8192$.
VI Conclusion
Analog computing is undergoing a revival, promising low-power and massively parallel computation capabilities for applications in signal processing and communications. In this paper, we introduce the concept of MiLAC as a generic analog computer that linearly operates with microwave signals, and analyze its application to signal processing for communications. A MiLAC can generally be modeled as a multiport microwave network made of tunable impedance (or admittance) components. At the ports of such a microwave network, the input signals are applied and the output signals are read. By using rigorous microwave theory, we derive the expression of the output of a MiLAC as a function of its input and the values of its admittance components. We show that a MiLAC can efficiently compute the LMMSE estimator in the analog domain, with significantly reduced computational complexity over conventional digital computing. In particular, it can be performed with complexity growing with the square of the input size, rather than the cube as with digital computing. This remarkable reduction in computational complexity is expected to play a key role in enabling beamforming and detection in gigantic MIMO communications.
Figure 3: Computational complexity of computing the LMMSE estimator given an observation vector with size $Y$, for various sizes $X$ of the unknown vector.
Appendix
In this appendix, we review the computational complexity of relevant matrix operations performed with classical digital computing. We define the computational complexity of an operation as the number of arithmetic real operations required to complete it, where these operations include addition, subtraction, multiplication, and division. Following this definition, we assume that computing the transpose or the conjugate transpose of a matrix requires no operations. For operations on complex numbers, we recall that addition and subtraction require two real operations, while multiplication requires six real operations.
VI-1 Scalar Product
Given two vectors $\mathbf{a}\in\mathbb{C}^{1\times N}$ and $\mathbf{b}\in\mathbb{C}^{N\times 1}$, the scalar product $\mathbf{a}\mathbf{b}$ requires $N-1$ complex additions and $N$ complex multiplications, since it is given by $[\mathbf{a}]_{1}[\mathbf{b}]_{1}+\ldots+[\mathbf{a}]_{N}[\mathbf{b}]_{N}$. Considering that complex additions and multiplications require two and six real operations, respectively, the complexity of the scalar product is approximately $8N$ real operations.
VI-2 Matrix-Vector Product
Given a matrix $\mathbf{A}\in\mathbb{C}^{M\times N}$ and a vector $\mathbf{b}\in\mathbb{C}^{N\times 1}$, the matrix-vector product $\mathbf{A}\mathbf{b}$ can be expressed as $\mathbf{A}\mathbf{b}=[\mathbf{a}_{1}\mathbf{b},\ldots,\mathbf{a}_{M}\mathbf{b}]^{T}$, where $\mathbf{a}_{m}\in\mathbb{C}^{1\times N}$ is the $m$th row of $\mathbf{A}$, for $m=1,\ldots,M$. Thus, it requires $M$ times the complexity of the scalar product $\mathbf{a}\mathbf{b}$, which is approximately $8MN$ real operations.
VI-3 Matrix-Matrix Product
Given two matrices $\mathbf{A}\in\mathbb{C}^{M\times N}$ and $\mathbf{B}\in\mathbb{C}^{N\times L}$, the matrix-matrix product $\mathbf{A}\mathbf{B}$ can be expressed as $\mathbf{A}\mathbf{B}=[\mathbf{A}\mathbf{b}_{1},\ldots,\mathbf{A}\mathbf{b}_{L}]$, where $\mathbf{b}_{\ell}\in\mathbb{C}^{N\times 1}$ is the $\ell$th column of $\mathbf{B}$, for $\ell=1,\ldots,L$. Thus, it requires $L$ times the complexity of the matrix-vector product $\mathbf{A}\mathbf{b}$, which is approximately $8LMN$ real operations.
VI-4 Matrix Inversion
Given an invertible matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, the algorithm commonly used to compute the matrix inverse $\mathbf{A}^{-1}$ is the Gauss’s method [22, Chapter 1]. Thus, we regard the complexity of matrix inversion as equal to the complexity of the Gauss’s method algorithm. It can be shown following [22, Chapter 1] that the Gauss’s method requires $N(N+1)/2$ divisions, $(2N^{3}+3N^{2}-5N)/6$ multiplications, and $(2N^{3}+3N^{2}-5N)/6$ subtractions. Consequently, the algorithm’s complexity is dominated by approximately $N^{3}/3$ multiplications and $N^{3}/3$ subtractions, which require a total of $38N^{3}/$ 3 real operations.